# Learning Degenerate Manifolds of Frustrated Magnets with Boltzmann Machines

### v1

@@ III ANNNI model
Spin configurations $\{\tilde{\bm{\sigma}}^{(s)}\}$, $s=1,2,\cdots,N_{\rm data}$, generated from the MC simulations are then used to train the RBM via the stochastic gradient descent update in Eq. (9). In particular, the data-dependent gradient term $\langle\partial E_{\bm{\theta}}/\partial\bm{\theta}\rangle_{\rm data}$ in evaluated by clamping the visible spins to the MC-sampled spin configurations. Because time-reversal symmetry remains unbroken at the multi-phase point, we set all visible and hidden biases to zero, leaving the coupling matrix $\mathbf{W}$ as the sole set of trainable parameters. To monitor training progress, we track the iteration-to-iteration change in $\mathbf{W}$, quantified by
@@
Fig. 3 plots $\varepsilon_{W}$ as a function of the iteration number $k$, providing a convenient diagnostic of convergence behavior. At early stages, $\varepsilon_{W}$ exhibits relatively large fluctuations, reflecting the rapid adjustment of the weights as the model aligns itself with the MC-sampled spin configurations. Within the first few tens of iterations, the amplitude of these fluctuations decreases markedly, indicating that the data-dependent and model-dependent contributions to the gradient are approaching balance. Beyond this transient regime, $\varepsilon_{W}$ decays smoothly toward zero over hundreds of iterations, eventually reaching a small, nearly constant plateau. This saturation signals that the RBM parameters have entered a steady-learning regime in which subsequent updates become minimal, consistent with convergence of the stochastic gradient descent procedure.
@@ IV.1 Ice-I phase
Figure 7: Comparison of the spin-spin correlation function $C(r)$ obtained from MC sampling and from RBM-generated configurations for the $T=0$ ice-I manifold of $\mathcal{H}_{\scriptsize\mbox{ice-I}}$. The inset shows $|C(r)|$ on a semi-logarithmic scale, highlighting the short-range nature of the correlations. The dashed line indicates an exponential fit, $\exp(-r/\ell)$, with correlation length $\ell\approx 0.95$.
The training data for the kagome ice-I phase were generated using standard Markov-chain Monte Carlo (MC) simulations on a $15\times 15$ lattice (with $N=3\times 15^{2}$ spins) at $T=0$. In contrast to the more constrained pyrochlore ice or the charge-ordered ice-II manifold discussed below, the kagome ice-I phase possesses an extensive set of local zero modes: many single-spin flips preserve the 2-in/1-out (or 1-in/2-out) constraint on every triangle. As a result, single-spin Metropolis updates efficiently explore the ice-I manifold without becoming trapped in restricted sectors of configuration space. As shown in Fig. 7, an RBM trained on these MC samples accurately reproduces the spin–spin correlation function, indicating that it has learned the essential statistics of the ice-I phase.
The correlation function itself reflects the characteristic short-range order imposed by the ice rules. The oscillatory behavior with a period of one lattice spacing arises from the underlying nearest-neighbor antiferromagnetic interaction of the Ising spins, while the strong overall exponential envelope, $e^{-r/\ell}$, evidences the geometric frustration that prevents long-range order. This exponential decay is clearly visible in the semi-logarithmic inset of Fig. 7, with a correlation length $\ell\sim\mathcal{O}(1)$ lattice spacing. The agreement between RBM-generated data (red circles) and MC results (blue triangles) across all distances demonstrates that the trained RBM successfully captures both the local ice-rule constraints and the rapidly decaying correlations characteristic of the kagome ice-I manifold.
@@ IV.2 Ice-II phase
To train RBMs for the ice-II phase, we performed MC simulations of the Hamiltonian in Eq. (15) at temperatures $T\lesssim K$, where the short-range charge interaction stabilizes staggered charge order. Throughout, the temperature $T$ is measured in units of the nearest-neighbor coupling $J$, and we set the charge–charge repulsion to $K=0.1J$. This order imposes additional constraints beyond the ice rules and greatly restricts the accessible configuration space. In a charge-ordered background, a single-spin flip almost always disrupts the $Q_{\alpha}=\pm 1$ pattern, creating either costly $Q=\pm 3$ monopoles or the less severe but still unfavorable misplaced $Q=\pm 1$ defects. As a result, naive single-spin Metropolis updates become inefficient. To ensure adequate sampling, our MC simulations therefore combine local single-spin flips with loop updates that flip spins along closed paths while preserving the underlying charge pattern.
The emergence of charge order also breaks a $Z_{2}$ symmetry associated with time-reversal. This implies that the effective RBM Hamiltonian in Eq. (7) must explicitly break time-reversal symmetry, highlighting the importance of including local bias fields. Indeed, unlike in the ice-I case, RBMs trained without local fields perform poorly for ice-II configurations. Therefore, in the RBM training discussed below, we treat $\bm{b}$, $\bm{b}^{\prime}$, and $\bm{W}$ as full sets of trainable parameters. These parameters are adjusted through the training process such that the marginal distribution Eq. (5) accurately approximates the Boltzmann distribution of the ice-II Hamiltonian at the given temperature, i.e. $\pi_{\bm{\theta}}(\bm{\sigma})\approx\pi_{\scriptsize\mbox{ice-II}}(\bm{\sigma};\,T)\propto\exp(-\mathcal{H}_{\scriptsize\mbox{ice-II}}/T)$.
Figure 8: Comparison of the spin-spin correlation function $C(r)$ obtained from Markov-chain MC simulations and RBM-generated configurations for the kagome ice-II phase at $T=K=0.1J$. The inset shows the absolute value $|C(r)|$ on a log-log scale, where the dashed line indicates a $1/r^{2}$ power-law decay.
@@
Figure 9: Histogram of RBM weights $W_{ij}$ for ice-I (red) and ice-II (blue) phases. The ice-I distribution is broad and symmetric, reflecting the highly fluctuating ice manifold, whereas the ice-II weights are sharply peaked near zero, consistent with the reduced configurational variability in the charge-ordered phase.
The histogram of trained RBM weights provides a useful window into how the model encodes the distinct correlation structures of kagome ice-I and ice-II. As shown in Fig. 9, the weight distribution for the ice-I manifold (red) is broad and nearly symmetric, with substantial support for both positive and negative couplings. This reflects the fluctuating character of ice-I: the two-in/one-out constraint enforces only short-range correlations and leaves a macroscopically degenerate configuration space, which the RBM captures through a diffuse set of visible-hidden couplings.
By contrast, the distribution learned for the ice-II phase (blue) is significantly narrower and more sharply peaked near zero. Long-range charge order reduces the variability of local spin environments, allowing the RBM to encode the data using smaller-magnitude couplings. The pronounced narrowing of the weight histogram thus mirrors the reduced configurational entropy and enhanced regularity of the charge-ordered phase. These weight statistics therefore offer a direct diagnostic of the underlying correlation hierarchy, naturally distinguishing the extensively degenerate ice-I manifold from the partially ordered ice-II phase.
@@
Figure 10: Histograms of the RBM local fields for the visible (light curves) and hidden (dark curves) layers at two temperatures within the ice-II phase: (a) $T=1.0$ and (b) $T=0.1$, in units of the nearest-neighbor AF coupling $J$ with charge–charge repulsion fixed at $K=0.1J$. At both temperatures, the visible-layer fields cluster around a single positive sign, while the hidden-layer fields peak around the opposite sign, consistent with the broken time-reversal symmetry of the charge-ordered ice-II phase.
The distributions of the local bias fields learned by the RBM provide a direct probe of how the model captures the broken symmetries and correlation structure of the ice-II phase. As shown in Fig. 10 for two temperatures within the ice-II regime, the visible-layer fields $\bm{b}$ (light curves) are consistently concentrated on the same sign, whereas the hidden-layer fields $\bm{b}^{\prime}$ form a distribution centered at the opposite sign. This clear sign separation persists at both $T=1.0$ and $T=0.1$, indicating that it is a robust feature of the trained models rather than a temperature-dependent artifact. The magnitude of the fields narrows as temperature decreases, reflecting the reduced configurational fluctuations deep in the charge-ordered phase, but the relative sign structure remains unchanged.
@@
The emergence of uniformly signed visible-layer fields is particularly significant. In the RBM effective energy, these fields act as local “magnetic fields” that bias the visible spins toward one orientation. A uniform sign in the visible-layer fields acts as an effective global bias on the physical spins and directly reflects the broken time-reversal $Z_{2}$ symmetry of the ice-II phase. Once monopole charges order, the local spin environments acquire a preferred orientation, and the RBM encodes this by assigning visible biases of consistent sign. The hidden-layer fields show the complementary structure: they peak around the opposite sign, indicating that the hidden units activate in a way that reinforces the broken-symmetry pattern. This opposing-sign structure in the two layers demonstrates that the RBM internalizes the essential symmetry breaking of ice-II and represents it through a coherent organization of its bias fields.

### v2

@@ III ANNNI model
Spin configurations ${\bm{\sigma}}^{(s)}$}, $s=1,2,\cdots,N_{\rm data}$, generated from Monte Carlo simulations are used to train the RBM via the stochastic gradient descent update in Eq. (9). The data-dependent gradient term $\langle\partial E_{\bm{\theta}}/\partial\bm{\theta}\rangle_{\rm data}$ is evaluated by clamping the visible spins to the MC-sampled configurations and computing the corresponding hidden-unit expectations from the exact conditional distribution. The model-dependent expectation is approximated using the standard contrastive divergence algorithm (CD-$k$), in which short Gibbs chains initialized from data configurations are evolved for a small number of steps to estimate the model averages. For the relatively simple ANNNI model considered here—where the configuration space is less degenerate and mixing is efficient—CD-$k$ provides a sufficiently accurate and stable estimate of the gradient.
Because time-reversal symmetry remains unbroken at the multiphase point, we fix all visible and hidden biases to zero, leaving the coupling matrix $\bm{W}$ as the only set of trainable parameters. To monitor the training progress, we track the iteration-to-iteration change in $\bm{W}$, quantified by
@@
Figure 3 shows $\varepsilon_{W}$ as a function of the iteration number $r$, providing a convenient diagnostic of convergence. At early stages, $\varepsilon_{W}$ exhibits relatively large fluctuations, reflecting the rapid reorganization of the weights as the model adjusts to the statistical structure of the MC-sampled configurations. Within the first few tens of iterations, the amplitude of these fluctuations decreases substantially, indicating that the data-dependent and model-dependent contributions to the gradient are approaching balance. Beyond this transient regime, $\varepsilon_{W}$ decays smoothly over hundreds of iterations and eventually reaches a small, nearly constant plateau. This saturation signals that the RBM parameters have entered a steady-learning regime in which subsequent updates are minimal, consistent with convergence of the stochastic gradient descent dynamics.
@@ IV.1 Ice-I phase
Figure 7: Comparison of the spin-spin correlation function $C(r)$ obtained from MC sampling and from RBM-generated configurations for the $T=0$ ice-I manifold of $\mathcal{H}_{\scriptsize\mbox{ice-I}}$. The inset shows $|C(r)|$ on a semi-logarithmic scale, highlighting the short-range nature of the correlations. The dashed line indicates an exponential fit, $\exp(-r/\ell)$, with correlation length $\ell\approx 0.89$.
The training data for the kagome ice-I phase were generated using standard Markov-chain Monte Carlo (MC) simulations on a $32\times 32$ lattice ($N=3\times 32^{2}$ spins) at $T=0$. Unlike the more constrained pyrochlore ice or the charge-ordered ice-II manifold discussed below, kagome ice-I possesses an extensive manifold of local zero modes: many single-spin flips preserve the 2-in/1-out (or 1-in/2-out) constraint on every triangle. Consequently, simple single-spin Metropolis updates efficiently explore the manifold without becoming confined to restricted sectors of configuration space.
For the RBM training, however, accurate sampling of the model distribution is essential in order to avoid systematic bias in the gradient estimation. As discussed in more detail in Appendix A, we therefore employ persistent contrastive divergence (PCD) rather than the standard CD-$k$ algorithm. In large and highly degenerate systems such as kagome ice, CD-kk—which repeatedly initializes short Gibbs chains from data configurations—can produce biased estimates of the model expectation values, potentially leading to spurious symmetry breaking or imbalance between degenerate sectors of the manifold. By contrast, PCD maintains a set of persistent Markov chains throughout training, allowing them to track the slowly evolving stationary distribution of the RBM. This substantially reduces gradient bias and ensures that the learned model faithfully represents the uniform $T=0$ ice manifold without artificial sector selection.
Figure 7 shows that the RBM trained on these MC samples quantitatively reproduces the spin–spin correlation function. The correlations display the characteristic short-range structure imposed by the ice rules: an oscillatory component with a period of one lattice spacing, originating from the nearest-neighbor antiferromagnetic interaction, modulated by a strong exponential envelope $e^{-r/\ell}$. The exponential decay, clearly visible in the semi-logarithmic inset, reflects geometric frustration that suppresses long-range order and yields a correlation length $\ell\sim\mathcal{O}(1)$ lattice spacing. Small deviations at large distances arise from finite-size effects. The close agreement between RBM-generated configurations (red circles) and MC data (blue triangles) across all distances demonstrates that the trained RBM faithfully captures both the local ice-rule constraints and the short-range correlated nature of the kagome ice-I manifold.
@@ IV.2 Ice-II phase
Figure 8: Comparison of the spin–spin correlation function $C(r)$ for the kagome ice-II phase at $T=0$ with nearest-neighbor charge repulsion $K=0.1J$. Results from Markov-chain Monte Carlo simulations are compared with correlations computed from RBM-generated configurations, demonstrating excellent agreement across all distances. The inset shows the absolute value $|C(r)|$ on a log–log scale; the dashed line indicates a $1/r^{2}$ power-law decay, highlighting the algebraic nature of correlations in the charge-ordered ice-II phase.
To train RBMs for the ice-II phase, we perform Monte Carlo simulations of the Hamiltonian in Eq. (15) at zero temperature, where the short-range charge interaction stabilizes a staggered charge-ordered state. Throughout, the nearest-neighbor charge repulsion is fixed at $K=0.1J$. The emergence of charge order imposes constraints beyond the conventional ice rules, substantially reducing the accessible configuration space and rendering the ground-state manifold far more restricted than in the ice-I case. In this charge-ordered background, a single-spin flip generically violates the local $Q_{\alpha}=\pm 1$ condition, either creating high-energy $Q=\pm 3$ monopole defects or producing misplaced $Q=\pm 1$ charges that disrupt the staggered arrangement. At $T=0$, where any defect carries a finite energy penalty, such local moves are strictly forbidden, and naive single-spin Metropolis updates become entirely ineffective.
The Monte Carlo sampling of the ice-II manifold at zero temperature is therefore carried out exclusively using nonlocal loop updates. These updates construct closed loops of spins whose collective reversal preserves the $Q_{\alpha}=\pm 1$ constraint on every triangle while maintaining the global staggered charge pattern. Because the net charge on each triangle remains unchanged, flipping an entire loop connects two distinct configurations within the same charge-ordered manifold without generating monopoles or other defects. This loop algorithm ensures ergodic exploration of the restricted ground-state manifold while exactly respecting the zero-energy constraints imposed by the Hamiltonian.
The onset of charge order also breaks a $Z_{2}$ symmetry associated with time-reversal. Consequently, the effective RBM Hamiltonian in Eq. (7) must be capable of explicitly breaking time-reversal symmetry, which underscores the necessity of including local bias fields in the model. In contrast to the ice-I case—where symmetry considerations allow the biases to be fixed to zero—RBMs trained without local fields fail to accurately represent ice-II configurations. Accordingly, in the RBM training for the ice-II phase, we treat $\bm{b}$, $\bm{b}^{\prime}$, and $\bm{W}$ as fully trainable parameters. These parameters are optimized such that the marginal distribution in Eq. (5) faithfully approximates the degenerate $T=0$ manifold, i.e., $\pi_{\bm{\theta}}(\bm{\sigma})\approx\pi_{\scriptsize\mbox{ice-II}}(\bm{\sigma};\,T=0)$, which is again a uniform distribution over all spin configurations satisfying the restricted ice rules: $\pi_{\scriptsize\mbox{ice-II}}(\bm{\sigma};\,T=0)=1/\mathcal{Z}$ Here $\mathcal{Z}=\exp(NS_{0})$ denotes the ground-state degeneracy, with residual entropy per spin $S_{0}\approx 0.108$.
@@
Figure 9: Histogram of the learned RBM couplings $W_{ij}$ for the kagome ice-I and ice-II phases. The ice-I distribution (blue) is relatively narrow and centered near zero, consistent with the highly degenerate and weakly constrained ice-I manifold. By contrast, the ice-II distribution (red) is noticeably broader, with larger-magnitude weights. This increased variance reflects the stronger constraints and reduced entropy of the charge-ordered ice-II phase.
The broader distribution of the learned couplings $\bm{W}$ in the ice-II phase can be understood as a consequence of its more strongly constrained and lower-entropy ground-state manifold. Compared with ice-I, which possesses an extensive set of local zero modes and a relatively flat, highly degenerate configuration space, the charge-ordered ice-II phase occupies a smaller and more structured subset of spin configurations. In particular, the additional charge-order constraint reduces the residual entropy and imposes more rigid correlations between spins. As shown in Fig. 9, the histogram of $W_{ij}$ for ice-II exhibits a visibly larger variance and heavier tails than that of ice-I, indicating that the RBM develops stronger visible–hidden couplings in order to encode these enhanced constraints. To faithfully represent the restricted ice-II manifold, the model must more sharply suppress configurations that violate the charge-order condition while enhancing the weight of symmetry-compatible states, which naturally leads to larger-magnitude parameters.
From the perspective of the effective RBM energy landscape, ice-I corresponds to a comparatively shallow and nearly isotropic manifold, reflected in the narrower and more concentrated weight distribution seen in Fig. 9. By contrast, the ice-II phase exhibits more pronounced energy basins associated with broken time-reversal symmetry and emergent sublattice order. Capturing these sharper distinctions requires stronger correlations mediated by the hidden units, resulting in the broader spread of $W_{ij}$ observed in the figure. In this sense, the enhanced variance of the weight distribution for ice-II directly mirrors its reduced entropy and increased structural coherence, and thus provides an informative diagnostic of how the RBM internalizes the underlying physical constraints.
@@
Figure 10: Histograms of the RBM local fields for the ice-II phase. Light curves correspond to the visible-layer biases, while dark curves denote the hidden-layer biases. The visible-layer fields predominantly cluster around one positive sign, whereas the hidden-layer fields peak around the opposite sign. This clear sign asymmetry reflects the spontaneous breaking of time-reversal symmetry associated with the charge-ordered ice-II phase.
@@
The distributions of the local bias fields learned by the RBM provide a direct probe of how the model captures the broken symmetries and correlation structure of the ice-II phase. As shown in Fig. 10, the visible-layer fields $\bm{b}$ (light curves) are consistently concentrated on the same sign, whereas the hidden-layer fields $\bm{b}^{\prime}$ form a distribution centered at the opposite sign. The emergence of uniformly signed visible-layer fields is particularly significant. In the RBM effective energy, these fields act as local “magnetic fields” that bias the visible spins toward one orientation. A uniform sign in the visible-layer fields acts as an effective global bias on the physical spins and directly reflects the broken time-reversal $Z_{2}$ symmetry of the ice-II phase. Once monopole charges order, the local spin environments acquire a preferred orientation, and the RBM encodes this by assigning visible biases of consistent sign. The hidden-layer fields show the complementary structure: they peak around the opposite sign, indicating that the hidden units activate in a way that reinforces the broken-symmetry pattern. This opposing-sign structure in the two layers demonstrates that the RBM internalizes the essential symmetry breaking of ice-II and represents it through a coherent organization of its bias fields.
@@ Acknowledgements.
Appendix A Numerical Implementation Details
As discussed in Sec. II, training the RBM amounts to learning a parametric probability distribution $\pi_{\bm{\theta}}(\bm{\sigma})$ that approximates the physical Boltzmann distribution $\pi_{\mathrm{phys}}(\bm{\sigma})$ defined in Eq. (8). Here $\bm{\sigma}=\{\sigma_{1},\dots,\sigma_{N}\}$ denotes a visible-layer spin configuration. The discrepancy between the two distributions is quantified by the Kullback–Leibler (KL) divergence,
$D_{\mathrm{KL}}\!\left(\pi_{\mathrm{phys}}\,\middle\|\,\pi_{\bm{\theta}}\right)=\sum_{\bm{\sigma}}\pi_{\mathrm{phys}}(\bm{\sigma})\ln\frac{\pi_{\mathrm{phys}}(\bm{\sigma})}{\pi_{\bm{\theta}}(\bm{\sigma})}.$ (16)
which measures how well the RBM reproduces the statistical weights of physical configurations. Minimizing $D_{\mathrm{KL}}$ with respect to the model parameters $\bm{\theta}=\{\bm{W},\bm{b},\bm{b}^{\prime}\}$ is equivalent to maximizing the likelihood of the training data under the RBM ansatz and constitutes the central optimization task. The parameters are initialized as $\bm{\theta}^{(0)}=\{\bm{W}^{(0)},\bm{b}^{(0)},\bm{b}^{\prime(0)}\}$, with vanishing biases $\bm{b}^{(0)}=\bm{b}^{\prime(0)}=0$ and weights $W_{ij}^{(0)}$ drawn independently from a normal distribution with zero mean and standard deviation $0.01$. Such small random initialization places the model near a high-entropy, weakly correlated state, allowing nontrivial correlations and symmetry breaking to emerge gradually during training.
Parameter updates follow stochastic gradient descent,
$\Delta\bm{\theta}^{(n)}=\bm{\theta}^{(n+1)}-\bm{\theta}^{(n)}=-\eta\,\nabla_{\bm{\theta}}D_{\mathrm{KL}},$ (17)
where $\eta$ denotes the learning rate. Differentiating the KL divergence yields, for example, for the weight matrix,
$\frac{\partial D_{\mathrm{KL}}}{\partial W_{ij}}=-\bigl\langle\sigma_{i}\tau_{j}\bigr\rangle_{\pi_{\mathrm{phys}}(\bm{\tau}\mid\bm{\sigma})}+\bigl\langle\sigma_{i}\tau_{j}\bigr\rangle_{\pi_{\bm{\theta}}(\bm{\sigma},\bm{\tau})}.$ (18)
This expression has a clear physical interpretation. The first term (“data average”) measures correlations between visible and hidden units when the visible configuration is clamped to a Monte Carlo sample drawn from the physical distribution. Because the RBM factorizes conditionally, the hidden expectation values can be computed exactly from $\pi_{\bm{\theta}}(\bm{\tau}\mid\bm{\sigma})$. The second term (“model average”) requires sampling from the full joint model distribution $\pi_{\bm{\theta}}(\bm{\sigma},\bm{\tau})$. This step is computationally challenging because evaluating exact expectations would require summing over an exponentially large configuration space through the partition function $Z_{\bm{\theta}}$. In practice, one therefore relies on Markov-chain Monte Carlo sampling within the RBM itself.
Because exact evaluation would require computing the partition function $Z_{\bm{\theta}}$ over an exponentially large configuration space, one resorts to Markov-chain Monte Carlo sampling within the RBM. A natural choice is Gibbs sampling, which alternately updates visible and hidden variables according to their exact conditional distributions:
$\bm{\sigma}^{(0)}\to\bm{\tau}^{(0)}\to\bm{\sigma}^{(1)}\to\bm{\tau}^{(1)}\to\cdots\to\bm{\sigma}^{(k)}\to\bm{\tau}^{(k)},$
where each transition is performed using the conditional probabilities $\pi_{\bm{\theta}}(\bm{\tau}\mid\bm{\sigma})$ and $\pi_{\bm{\theta}}(\bm{\sigma}\mid\bm{\tau})$. Because each conditional update satisfies detailed balance, this Markov chain converges, in the limit of long sampling time, to the stationary Boltzmann distribution of the RBM. In principle, one could run this chain to equilibrium at every parameter update and use the resulting samples to estimate the model average in Eq. (18). However, such full equilibration at every training step is computationally prohibitive for the large spin systems considered here.
In principle, one could equilibrate the Gibbs chain at each parameter update to obtain unbiased estimates of the model expectation. In practice, however, full equilibration is computationally prohibitive for large systems such as the kagome lattices studied here. Contrastive divergence (CD-$k$) addresses this issue by initializing the Gibbs chain from a configuration drawn from the training dataset and performing only a small number $k$ of alternating updates before estimating the model expectation. While CD-$k$ often performs surprisingly well for moderate system sizes, it provides a biased gradient estimate because the truncated chain samples from a distribution that remains closer to the data distribution than to the true model distribution [87].
This bias becomes particularly pronounced in systems with large configuration spaces, slow mixing, or highly degenerate manifolds. In the kagome ice-I phase, for example, the ground-state manifold consists of symmetry-related sectors with equal statistical weight. Physical observables such as the unit-cell spin sum $Q_{\alpha}={-1,1}$ must therefore exhibit symmetric distributions. During CD-$k$ training, however, we observe that the model gradually develops a preference for one sector. This spurious symmetry breaking originates from insufficient equilibration of the Gibbs chain: because each chain is repeatedly reinitialized from the data and evolved only briefly, it fails to sample the full model distribution and instead drifts toward metastable regions of configuration space. The resulting gradient estimate is systematically biased.
To reduce this bias, we employ persistent contrastive divergence (PCD) [88]. The essential modification relative to CD-$k$ is that the Markov chains used to estimate the model expectation are not reinitialized from the data at every iteration. Instead, a set of persistent chains is maintained throughout the entire training process. At each parameter update, these chains are advanced by a small number of Gibbs steps, and their current states are used to estimate the model expectation. The conceptual advantage of PCD relies on the fact that the model parameters evolve gradually during training, provided the learning rate is sufficiently small. Consequently, the stationary distribution $\pi_{\bm{\theta}}$ changes only adiabatically between successive updates. A persistent Markov chain that was approximately equilibrated with respect to $\pi_{\bm{\theta}^{(n)}}$ remains close to equilibrium for $\pi_{\bm{\theta}^{(n+1)}}$, requiring only a few additional Gibbs steps to re-equilibrate. In this sense, PCD tracks the slowly drifting Boltzmann distribution of the RBM, rather than repeatedly restarting the sampling process from the data manifold.
From a statistical perspective, CD-$k$ estimates the model expectation using a distribution obtained after $k$ applications of the Gibbs transition operator to the data distribution, whereas PCD estimates it using a distribution that remains close to the stationary model distribution throughout training. In the limit of vanishing learning rate and sufficiently long Gibbs chains, PCD approaches exact maximum-likelihood learning. In practice, even a small number of Gibbs steps per update significantly reduces the systematic bias compared to CD-$k$. In our simulations of kagome spin ice, PCD eliminates the artificial imbalance between symmetry-related sectors observed under CD-$k$ and restores the expected symmetry of observables within the degenerate manifold. All results presented for both ice-I and ice-II phases are therefore obtained using the PCD algorithm.
To assess the RBM’s capacity to represent these extensively degenerate manifolds, we consider systems of size $N=3\times 32^{2}$ (ice-I) and $N=3\times 24^{2}$ (ice-II). The training dataset consists of 5000 independent Monte Carlo configurations, generated with sufficient thinning to suppress autocorrelation effects. We employ a hidden layer with $N_{h}=5000$ units, providing substantial representational capacity while maintaining stable optimization. For the ice-I phase, the visible and hidden biases $\bm{b}$ and $\bm{b}^{\prime}$ are fixed to zero in order to preserve the time-reversal symmetry of the Hamiltonian. In contrast, nonzero local fields are allowed for ice-II as time-reversal symmetry is broken in the presence of magnetic charge order. Training is performed for 2000 epochs using the Adam optimizer, with a scheduled learning rate: $\eta=10^{-4}$ for the first 1000 epochs and $\eta=10^{-5}$ for the remaining 1000 epochs. No weight decay is applied. All computations are carried out in PyTorch with GPU acceleration on an NVIDIA RTX A6000.